10. Determine the probability density function for the cumulative distribution function



shown below.




F {0, x < −2



0.25x + 0.5, −2 ≤ x < 1



0.5x + 0.25, 1 ≤ x < 1.5



1 x > 1.5




Determine also the following probabilities.



(a) P(X > −1)



(b)P(X < 1.3)



(c) P(−1.5 ≤ X ≤ 1.8)



(d)P(X < 1.5)

1
Expert's answer
2022-03-07T17:14:04-0500

f(x) - density function, then f(x)=F'(x), so

"f(x)=(0.25x+0.5)'=0.25, -2\u2264x<1"

"f(x)=(0.5x+0.25)'=0.5, 1\u2264x<1.5"

"f(x)=0," otherwise

a) "P(X>-1)=1-P(X\u2264-1)=1-F(-1)=1-0.25*(-1)-0.5=0.75"

b) "P(X<1.3)=F(1.3)=0.5*1.3+0.25=0.9"

c) "P(-1.5\u2264X\u22641.8)=F(1.8)-F(-1.5)=1-0.25*(-1.5)-0.5=0.875"

d) "P(x<1.5)=F(1.5)=1"